OpenStudy (anonymous):

Solve the following system of equations using matrices. 8x + 7y = 5 4x – 9y = 65 Answer (5, -5) (5, 65) (7, -7) (65, 5)

OpenStudy (anonymous):

@Chubba can you help me with this one?

OpenStudy (anonymous):

whoa nelly \[8x+7y=5\\ 4x-9y=65\] you need the method or the answer?

OpenStudy (anonymous):

if you just want the answer, it is in this link

OpenStudy (anonymous):

I would like the method, in order to solve it in the future. It confused me because of the matrices

OpenStudy (anonymous):

matrix method is a pain in the neck, and worse to write multiply the first equation by 9, the second one by 7, then add them up the y terms will drop out

OpenStudy (anonymous):

or if you like you can check which of the 4 solutions you are given work in both equations that actually takes longer to do

OpenStudy (anonymous):

So you can solve this in other ways than the question stated?

OpenStudy (anonymous):

\[8x+7y=5\\ 4x-9y=65\] \[9(8x+7y=5)\\ 7(4x-9y=65)\]

OpenStudy (anonymous):

oh hell yes

OpenStudy (anonymous):

Okay, I wasnt sure.

OpenStudy (anonymous):

\[72x+63y=45\\ 28x-63y=455\]

OpenStudy (anonymous):

add and get \(100x=500\) making \(x=5\)

OpenStudy (anonymous):

unless you have to do this in an in class exam, all of this is donkey work why they invented computers i would just use wolfram and get it over with in two seconds

OpenStudy (anonymous):

sorry in algebra right now myself. and am sitting in class right now. my method would have been plugging the ordered pairs in to see which one was true. then go from there.

OpenStudy (anonymous):

okay thanks I didn't know about that site.

OpenStudy (anonymous):

i sent the link, note the syntax i wrote

OpenStudy (anonymous):

and i would do that as long as there is multiple options to chose from, otherwise you should know how to work the method out.